With the circuit shown, if the ac voltage source has a constant voltage but increases in frequency what happens to magnitudeVs R Choose... R Vs Choose... c R WW Vs Choose... R
Answer #1

Ті eller mm L impedance, Z = R + JWL magnitude IZL) = √ R² + W²22 i= IVI √ R² + W²22 If frequency, w increases i decreases 2Iri impedance, Z= RIIR mann www = R.R - ald Vs R&R Current , i=1 Vsi So, if w- i. 211) R/2 R increases Remains same as does nS Ini ER Vs impedance, z= Rt Xc = R-- WE magnitude, 1Z1 = R² + 1 Lok (2 . Current, IVI 121 IV R²+ 2 w²c2 magnitude, Izl= R²+